Session: Oct 2010 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Sun Oct 31 13:38:55 2010) Success:

Yebes (Paco Colomer at Sun Oct 31 14:53:28 2010) Success:
Sync problem at 05:43:50 UT (during 1 minute).

Onsala (Michael Lindqvist at Mon Nov 01 05:26:26 2010) Success:
No known problems.

Effelsberg (Uwe Bach at Tue Nov 02 08:46:46 2010) Success:

Metsahovi (Minttu Uunila at Thu Nov 04 09:15:03 2010) Minor failures:
BBC01 is suspected to be broken although reports 'lock'.

Metsahovi (Minttu Uunila at Thu Nov 04 09:39:18 2010) Minor failures:
BBC02 probably unlocked, didn't want to lock to the new low astro K band continuum setup frequency.

Robledo (Isotuela at Wed Nov 10 16:17:54 2010) Minor failures: TSYS WIND
Diode failure. Zenith calibration and ambient load instead.